OpenStudy (anonymous):

You weigh out a sample of 2.45 grams of fluorine. How many atoms are in your sample? I think I have an idea of how to get it but I wanna be sure

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

calculate the number of moles of Fluorine atom then multiply by 6.02 x 10 ^23

OpenStudy (anonymous):

ok so 2.45/18.99 g/mol=0.129 mol * (6.02*10^23)

OpenStudy (anonymous):

0.129 x 2 x 6.02 x 10^23

OpenStudy (anonymous):

ok and is 7.77*10^22 atoms of F correct?

OpenStudy (anonymous):

where did the 2 in your equation come from @Ezizio

OpenStudy (anonymous):

what u just calculated was the number of F2 molecules

OpenStudy (anonymous):

they are asking for the number of F atoms

OpenStudy (anonymous):

yes and ! divided yours by 2 and got 7.77*10^22 atoms of F

OpenStudy (anonymous):

you need to multiply by 2

OpenStudy (anonymous):

2.45 / 18.99 = moles of F2 molecules

OpenStudy (anonymous):

(2.45 / 18.99) x 2 = moles of F atoms

OpenStudy (anonymous):

yup, and 2.45 / 18.99 = moles of F2 molecules=0.129 mol 0.129 mol* 6.02*10^23=7.77* 10^22 atoms of F

OpenStudy (anonymous):

april sages, what you've just calculated is the number of F2 molecules

OpenStudy (anonymous):

1 mole = 6.02 x 10^ 23 particles

OpenStudy (anonymous):

so you just found out 7.77 x 10^ 22 particles of F2

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

no atoms?

OpenStudy (anonymous):

they want atom, there is a difference between atom and molecules

OpenStudy (anonymous):

ok show me how to get atoms

OpenStudy (anonymous):

multiply your answer by 2

OpenStudy (anonymous):

so (7.77 x 10^ 22)*2

OpenStudy (anonymous):

yes, that is for number of F atoms

OpenStudy (anonymous):

so 1.554*10^23 is my final answer?

OpenStudy (anonymous):

@Ezizio

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!

OpenStudy (anonymous):

np